### v1

@@ 5 Ablations and Extra Invesitigations
5.1 Ablations on PASS
To evaluate the effectiveness of PASS, we pose two interesting questions about the design of its components: () how do visual prompts and model weights contribute? () is the recurrent mechanism crucial for mask finding? To answer the above questions, we conduct a series of ablation studies utilizing a pre-trained ResNet- on CIFAR-. The extensive investigations contain () dropping either the visual prompt or model weights; () destroy the recurrent nature in our hypernetwork, such as using a Convolutional Neural Network (CNN) or a Multilayer Perceptron (MLP) to replace LSTM. The results are collected in Table 3. We observe that ❶ The exclusion of either the visual prompt or model weights leads to a pronounced drop in test accuracy (e.g., and respectively at channel density), indicating the essential interplay role of both visual prompt and model weights in sparsification. ❷ If the recurrent nature in our design is destroyed, i.e., MLP or CNN methods variants, it suffers a performance decrement (e.g., and respectively at channel density). It implies a Markov property during the sparsification of two adjacent layers, which echoes the sparsity pathway findings in Wang et al. [40].
5.2 Ablations on Visual Prompt
@@
A visual prompt is a patch integrated with the input, as depicted in Figure 1. Two prevalent methods for incorporating the visual prompt into the input have been identified in the literature [35, 62]:() Adding to the input (abbreviated as “Additive visual prompt”. ()Expanding around the perimeter of the input, namely, the input is embedded into the central hollow section of the visual prompt (abbreviated as “Expansive visual prompt”). As discussed in section 5.1, visual prompt (VP) plays a key role in PASS. Therefore, we pose such a question:How do the strategies and size of VP influence the performance of PASS? To address this concern, we conduct experiments with “Additive visual prompt” and “Expansive visual prompt” respectively on CIFAR- using a pre-trained ResNet- under , and channel sparsities, and we also show the performance of PASS with varying the VP size from to . The results are shown in Figure 4. We conclude that ❶ “Additive visual prompt” performs better than “Expansive visual prompt” across different sparsities. The disparity might be from the fact that “Expansive visual prompt” requires resizing the input to a smaller dimension, potentially leading to information loss, a problem that “Additive visual prompt” does not face. ❷ The size of VP impacts the performance of PASS. We observe that test accuracy initially rises with the increase in VP size but starts to decline after reaching a peak at size . A potential explanation for this decline is that the larger additive VP might overlap a significant portion of the input, leading to the loss of crucial information.
5.3 Impact of Hidden Size in HyperNetwork
@@
It is well-known that model size is an important factor impacting its performance, inducing the question how does the size of hypernetwork influence the performance of PASS?. To address this concern, we explore the impact of the hypernetwork hidden sizes on PASS by varying the hidden size of the proposed hypernetwork from to and evaluate its performance on CIFAR using a pre-trained ResNet- model under , and channel sparsity respectively. The results are presented in the left figure of Figure 5. We observe that the hidden size of the hypernetwork doesn’t drastically affect the accuracy. While there are fluctuations, they are within a small range, suggesting that the hidden size is not a dominant factor in influencing the performance of PASS.
5.4 Uniform Pruning VS Global Pruning
@@
When converting the channel-wise importance scores into the channel masks, there are two prevalent strategies: () Uniform Pruning. [97, 81] It prunes the channels of each layer with the lowest important scores by the same proportion. () Global Pruning. [81, 20] It prunes channels with the lowest important scores from all layers, leading to varied sparsity across layers. In this section, we evaluate the performance of global pruning and uniform pruning for PASS on CIFAR- using a pre-trained ResNet-, with results presented in Figure 5. We observe that global pruning consistently yields higher test accuracy than uniform pruning, indicating its superior suitability for PASS, also reconfirming the importance of layer sparsity in sparsifying neural networks [98, 81].
@@
We present the channel sparsity learned by PASS with global pruning on CIFAR- and Tiny-ImageNet using a pre-trained ResNet- in Table 4 in Table 4. It can be observed that channel sparsity is generally higher in the top layers and lower in the bottom layers of the network.
Layer Fully Dense #Channels CIFAR-100 Tiny-ImageNet Sparsity Sparsity Sparsity Sparsity Layer 1 - conv1 Layer 2 - layer1.0.conv1 Layer 3 - layer1.0.conv2 Layer 4 - layer1.1.conv1 Layer 5 - layer1.1.conv2 Layer 6 - layer2.0.conv1 Layer 7 - layer2.0.conv2 Layer 8 - layer2.0.downsample.0 Layer 9 - layer2.1.conv1 Layer 10 - layer2.1.conv2 Layer 11 - layer3.0.conv1 Layer 12 - layer3.0.conv2 Layer 13 - layer3.0.downsample.0 Layer 14 - layer3.1.conv1 Layer 15 - layer3.1.conv2 Layer 16 - layer4.0.conv1 Layer 17 - layer4.0.conv2 Layer 18 - layer4.0.downsample.0 Layer 19 - layer4.1.conv1 Layer 20 - layer4.1.conv2 Layer 21 - Linear
6 Complexity Analysis of the Hypernetwork
@@
In this section, we provide a comprehensive analysis about the complexity of the Hypernetwork. Regarding the impact on time complexity, our recurrent hyper-network is designed for efficiency. The channel masks are pre-calculated, eliminating the need for real-time generation during both the inference and subnetwork fine-tuning phases. Therefore, the recurrent hyper-network does not introduce any extra time complexity during the inference and the fintune-tuning phase. The additional computing time is limited to the phase of channel mask identification. Moreover, the hyper-network itself is designed to be lightweight. The number of parameters it contributes to the overall model is minimal, thus ensuring that any additional complexity during the mask-finding phase is negligible. This claim is substantiated by empirical observations: the hyper-network accounts for only about % to % of the total model parameters across various architectures such as ResNet-18/34 and VGG-16, as illustrated in Table 7. Additionally, we assessed the training time per epoch with and without the hyper-network during the channel mask identification phase. Our findings in Table 5 indicate that the inclusion of the LSTM network has a marginal effect on these durations, further affirming the efficiency of our approach.
@@
7 Conclusion
Appendix A Data Availability Statement
The datasets used in this paper are sourced from public repositories. The download link are shown in Table 6.
| Dataset | Download Links |
| ImageNet | https://image-net.org/download |
| TinyImageNet | http://cs231n.stanford.edu/tiny-imagenet-200.zip |
| CIFAR-10 | https://www.cs.toronto.edu/~kriz/cifar-10-python.tar.gz |
| CIFAR-100 | https://www.cs.toronto.edu/~kriz/cifar-100-python.tar.gz |
| DTD | https://www.robots.ox.ac.uk/~vgg/data/dtd/download/dtd-r1.0.1.tar.gz |
| StandfordCars | https://www.kaggle.com/datasets/jessicali9530/stanford-cars-dataset |
| Food101 | https://data.vision.ee.ethz.ch/cvl/datasets_extra/food-101/ |
Appendix B Parameters of Hypernetworks
| ResNet-18 (11M) | ResNet-34 (21M) | ResNet-50 (25M) | VGG-16 (138M) | |
Appendix C Implementation Details

### v2

@@ 5 Ablations and Extra Invesitigations
Ablations on PASS. To evaluate the effectiveness of PASS, we pose two interesting questions about the design of its components: () how do visual prompts and model weights contribute? () is the recurrent mechanism crucial for mask finding? To answer the above questions, we conduct a series of ablation studies utilizing a pre-trained ResNet- on CIFAR-. The extensive investigations contain () dropping either the visual prompt or model weights; () destroy the recurrent nature in our hypernetwork, such as using a Convolutional Neural Network (CNN) or a Multilayer Perceptron (MLP) to replace LSTM. The results are collected in Table 3. We observe that ❶ The exclusion of either the visual prompt or model weights leads to a pronounced drop in test accuracy (e.g., and respectively at channel density), indicating the essential interplay role of both visual prompt and model weights in sparsification. ❷ If the recurrent nature in our design is destroyed, i.e., MLP or CNN methods variants, it suffers a performance decrement (e.g., and respectively at channel density). It implies a Markov property during the sparsification of two adjacent layers, which echoes the sparsity pathway findings in [39].
@@
Ablations on Visual Prompt. A visual prompt is a patch integrated with the input, as depicted in Figure 1. Two prevalent methods for incorporating the visual prompt into the input have been identified in the literature [6, 1]:() Adding to the input (abbreviated as “Additive visual prompt”. ()Expanding around the perimeter of the input, namely, the input is embedded into the central hollow section of the visual prompt (abbreviated as “Expansive visual prompt”). As discussed in section 5, visual prompt (VP) plays a key role in PASS. Therefore, we pose such a question:How do the strategies and size of VP influence the performance of PASS? To address this concern, we conduct experiments with “Additive visual prompt” and “Expansive visual prompt” respectively on CIFAR- using a pre-trained ResNet- under , and channel sparsities, and we also show the performance of PASS with varying the VP size from to . The results are shown in Figure 4. We conclude that ❶ “Additive visual prompt” performs better than “Expansive visual prompt” across different sparsities. The disparity might be from the fact that “Expansive visual prompt” requires resizing the input to a smaller dimension, potentially leading to information loss, a problem that “Additive visual prompt” does not face. ❷ The size of VP impacts the performance of PASS. We observe that test accuracy initially rises with the increase in VP size but starts to decline after reaching a peak at size . A potential explanation for this decline is that the larger additive VP might overlap a significant portion of the input, leading to the loss of crucial information.
@@
Impact of Hidden Size in HyperNetwork It is well-known that model size is an important factor impacting its performance, inducing the question how does the size of hypernetwork influence the performance of PASS?. To address this concern, we explore the impact of the hypernetwork hidden sizes on PASS by varying the hidden size of the proposed hypernetwork from to and evaluate its performance on CIFAR using a pre-trained ResNet- model under , and channel sparsity respectively. The results are presented in the left figure of Figure 5. We observe that the hidden size doesn’t drastically affect the accuracy. While there are fluctuations, they are within a small range, suggesting that the hidden size is not a dominant factor in influencing the performance of PASS.
@@
Uniform Pruning VS Global Pruning. When converting the channel-wise importance scores into the channel masks, there are two prevalent strategies: () Uniform Pruning. [35, 18] It prunes the channels of each layer with the lowest important scores by the same proportion. () Global Pruning. [18, 13] It prunes channels with the lowest important scores from all layers, leading to varied sparsity across layers. In this section, we evaluate the performance of global pruning and uniform pruning for PASS on CIFAR- using a pre-trained ResNet-, with results presented in Figure 5. We observe that global pruning consistently yields higher test accuracy than uniform pruning, indicating its superior suitability for PASS, also reconfirming the importance of layer sparsity in sparsifying neural networks [29, 18].
6 Conclusion
Acknowledgements
The authors acknowledge the use of resources provided by the Isambard-AI National AI Research Resource (AIRR) and the Dutch national e-infrastructure, supported by the SURF Cooperative (Project EINF-17091). Isambard-AI is operated by the University of Bristol and funded by the UK Government’s Department for Science, Innovation and Technology (DSIT) via UK Research and Innovation and the Science and Technology Facilities Council [ST/AIRR/I-A-I/1023]. Finally, we thank the anonymous reviewers for their insightful comments, which significantly improved the quality of this paper.
Appendix A Parameters of Hypernetworks
| ResNet-18 (11M) | ResNet-34 (21M) | ResNet-50 (25M) | VGG-16 (138M) | |
Appendix B Implementation Details
Appendix C Learned Channel Sparsity
@@
We present the channel sparsity learned by PASS on CIFAR- and Tiny-ImageNet using a pre-trained ResNet- in Table 6. Our observations indicate that channel sparsity is generally higher in the top layers and lower in the bottom layers of the network.
| Layer | Fully Dense #Channels | CIFAR-100 | Tiny-ImageNet | ||
| Sparsity | Sparsity | Sparsity | Sparsity | ||
| Layer 1 - conv1 | |||||
| Layer 2 - layer1.0.conv1 | |||||
| Layer 3 - layer1.0.conv2 | |||||
| Layer 4 - layer1.1.conv1 | |||||
| Layer 5 - layer1.1.conv2 | |||||
| Layer 6 - layer2.0.conv1 | |||||
| Layer 7 - layer2.0.conv2 | |||||
| Layer 8 - layer2.0.downsample.0 | |||||
| Layer 9 - layer2.1.conv1 | |||||
| Layer 10 - layer2.1.conv2 | |||||
| Layer 11 - layer3.0.conv1 | |||||
| Layer 12 - layer3.0.conv2 | |||||
| Layer 13 - layer3.0.downsample.0 | |||||
| Layer 14 - layer3.1.conv1 | |||||
| Layer 15 - layer3.1.conv2 | |||||
| Layer 16 - layer4.0.conv1 | |||||
| Layer 17 - layer4.0.conv2 | |||||
| Layer 18 - layer4.0.downsample.0 | |||||
| Layer 19 - layer4.1.conv1 | |||||
| Layer 20 - layer4.1.conv2 | |||||
| Layer 21 - Linear | |||||
Appendix D Complexity Analysis of the Hypernetwork
@@
In this section, we provide a comprehensive analysis about the complexity of the Hypernetwork. (1) Regarding the impact on time complexity, our recurrent hyper-network is designed for efficiency. The channel masks are pre-calculated, eliminating the need for real-time generation during both the inference and subnetwork fine-tuning phases. Therefore, the recurrent hyper-network does not introduce any extra time complexity during the inference and the fintune-tuning phase. The additional computing time is limited to the phase of channel mask identification. (2) Moreover, the hyper-network itself is designed to be lightweight. The number of parameters it contributes to the overall model is minimal, thus ensuring that any additional complexity during the mask-finding phase is negligible. This claim is substantiated by empirical observations: the hyper-network accounts for only about % to % of the total model parameters across various architectures such as ResNet-18/34 and VGG-16, as illustrated in Table 4. (3) Additionally, we assessed the training time per epoch with and without the hyper-network during the channel mask identification phase. Our findings in Table 7 indicate that the inclusion of the LSTM network has a marginal effect on these durations, further affirming the efficiency of our approach.
@@
Appendix E Difference between our proposed PASS and dynamic neural network
There are two fundamental differences between our proposed PASS and dynamic neural network. (1) The hyper-network in our proposed PASS is not ‘dynamic’. Dynamic neural networks, as categorized in the literature, are networks capable of adapting their structures or parameters conditioned in a sample-dependent manner, as outlined in [14]. In contrast, the hyper-network within our PASS framework does not exhibit this ‘dynamic’ nature. It is designed to be dependent on a visual prompt (task-specific), as opposed to dynamically adjusting to input samples. This hyper-network’s role is confined to the channel mask identification phase and is not employed during the inference phase. Therefore, it is fundamentally different from dynamic neural networks. (2) Their goals are different. The fundamental goal of the hyper-network in PASS is distinct from that of dynamic neural networks. While the latter focuses on adapting their architecture or parameters based on input samples, our hyper-network is specifically engineered for the integration of visual prompts with the statics of model weights.